# Fabian Mauthe

### v1

@@ 4 Description of Data Sets
Prognosis
Link
26
NASA - Bearing Data Set
J. Lee, H. Qiu, G. Yu, J. Lin, and Rexnord Technical Services. IMS, University of Cincinnati. "Bearing Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Mechanical component
@@
Prognosis
Link
27
NASA - Capacitor Electrical Stress
J. Renwick, C. Kulkarni, and J Celaya "Capacitor Electrical Stress Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Electrolytic Capacitors
Prognosis
Link
28
NASA - Capacitor Electrical Stress 2
J. Celaya, C. Kulkarni, G. Biswas, and K. Goebel "Capacitor Electrical Stress Data Set - 2", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Electrolytic Capacitors
Prognosis
Link
29
NASA - CFRP Composites Data Set
Abhinav Saxena, Kai Goebel, Cecilia C. Larrosa, and Fu-Kuo Chang "CFRP Composites Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Materials
Carbon Fibre-Reinforced Polymer Composite
Prognosis
Link
30
NASA - HIRF Battery
C. Kulkarni, E. Hogge, C. Quach and K. Goebel "HIRF Battery Data Set", NASA Ames Prognostics Data Repository (http://ti.arc.nasa.gov/project/prognostic-data-repository), NASA Ames Research Center, Moffett Field, CA
Electrical component
@@
Prognosis
Link
31
NASA - IGBT
J. Celaya, Phil Wysocki, and K. Goebel "IGBT Accelerated Aging Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Transistor
Prognosis
Link
32
NASA - Li-ion Battery Aging Datasets
B. Saha and K. Goebel. "Battery Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
@@
Prognosis
Link
33
NASA - Milling Data set
A. Agogino and K. Goebel. BEST lab, UC Berkeley. "Milling Data Set ", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Manufacturing process
Milling
Progrognosis
Link
34
NASA - MOSFET Thermal Overstress Aging
J. R. Celaya, A. Saxena, S. Saha, and K. Goebel "MOSFET Thermal Overstress Aging Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Transistor
Prognosis
Link
35
NASA - Randomized Battery Usage Data Set
B. Bole, C. Kulkarni, and M. Daigle "Randomized Battery Usage Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
@@
Prognosis
Link
36
NASA - Small Satellite Power Simulation
C. Kulkarni and A. Guarneros "Small Satellite Power Simulation Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
@@
Diagnosis
Link
37
NASA - Turbofan engine degradation simulation data set
A. Saxena and K. Goebel. "Turbofan Engine Degradation Simulation Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Drive technology
Aircraft Engine*
Prognosis
Link
38
NASA - Turbofan engine degradation simulation data set 2
M. Chao, C.Kulkarni, K. Goebel and O. Fink (2021). "Aircraft Engine Run-to-Failure Dataset under real flight conditions", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Drive technology
Aircraft Engine*
Prognosis
Link
39
NIST - Robot Arm Position Accuracy
National Institute of Standards and Technology NIST (USA), Engineering Laboratory, Intelligent Systems Division
Robotic
Industrial Robot
Diagnosis
Link
40
OpenEI - Evaluation of Building Fault
Lin Guanjing and Robin Mitchel, Lawrence Berkeley National Laboratory, (2019)
Building
Building
Diagnosis
Link
41
OpenEI - Gearbox Fault Diagnosis
Yogesh Pandy, National Renewable Energy Laboratory
Mechanical component
Gear
Fault detection
Link
42
PHM Data Challenge 2008 - Turbofan
NASA Prognostics Center of Excellence
Drive technology
Aircraft Engine*
Prognosis
Link
43
PHM Data Challenge 2009 - Gearbox Fault Detection
PHM Society, Gearbox fault detection data set, 2010
Mechanical component
Gear System (Gears, Bearings, Shafts)
Diagnosis
Link
44
PHM Data Challenge 2010 - CNC milling machine cutters
X. Li, B.S. Lim, J.H. Zhou, S. Huang, S.J. Phua, K.C. Shaw, and M.J. Er; Singapore Institute of Manufacturing Technology, 71 Nanyang Drive, Singapore 638075 School of Electrical and Electronic Engineering, Nanyang Technological University, Nanyang Avenue, Singapore 639798
Manufacturing process
CNC Milling Machine Cutters
Prognosis
Link
45
PHM Data Challenge 2011 - Anemometer Fault Detection
Creators Unknown
Mechanical component
Anemometer
Fault detection
Link
46
PHM Data Challenge 2013 - Unknown
Creators Unknown
Unknown
Unknown
Health assessment
Link
47
PHM Data Challenge 2014 - Unknown
Creators Unknown
Unknown
Unknown
Health assessment
Link
48
PHM Data Challenge 2015 - Plant Fault Detection
Creators Unknown
Production system
Industrial Plant Monitoring
Diagnosis
Link
49
PHM Data Challenge 2016 - Semiconductor CMP
Crystec Technology Trading GmbH
Manufacturing process
Chemicalmechanical Planarization System
Prognosis
Link
50
PHM Data Challenge 2017 - Bogie Vehicle
Creators Unknown
Mechanical component
Bogie
Diagnosis
Link
51
PHM Data Challenge 2018 - Ion Mill in Waver Manufacturing
Kai Goebel (at this time at NASA)
Production system
Ion Mill Etching Tool
Prognosis
Link
52
PHM Data Challenge 2019 - Fatigue Cracks
Creators Unknown
Materials
Aluminum Structure with Dynamic Loading
Prognosis
Link
53
PHM Data Challenge 2020 Europe - Filtration System
Eker, O. F., Camci, F., & Jennions, I. K. (2016). Physics-based prognostic modelling of filter clogging phenomena. Mechanical Systems and Signal Processing, 75, pp. 395-412.
Process technology
Filtration
Prognosis
Link
54
PHM Data Challenge 2021 - Turbofan 2
M. Chao, C.Kulkarni, K. Goebel and O. Fink (2021). "Aircraft Engine Run-to-Failure Dataset under real flight conditions", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA (additional validation data compared to data set NASA - Turbofan engine degradation simulation data set 2)
Drive technology
Aircraft Engine*
Prognosis
Link
55
PHM Data Challenge 2021 Europe: SCARA-robot
Swiss Centre for Electronics and Microtechnology (CSEM)
Robotic
Industrial Robot
Diagnosis
Link
56
PHM Data Challenge 2022 - Rock Drills
Epiroc Rock Drills AB, Örebro, 702 25, Sweden and Linköping University, Linköping, 581 83, Sweden
Mechatronic system
Rock Drills
Diagnosis
Link
57
PHM Data Challenge 2022 Europe - PCB Production line
Danilo Giordano and Martino Trevisan, (2022). "PHME Data Challenge". European conference of the prognostics and health management society. / BITRON: Strada del Portone 95 - 10095 Grugliasco (TO), Italy, https://www.bitron.com/en
Production system
PCB Production Line
Diagnosis
Link
58
PHM Data Challenge 2023 - Gearbox
PHM Society, Gearbox data set, 2023
Mechanical component
Gear
Diagnosis
Link
59
PHM Data Challenge 2023 Asia Pacific - Experimental Propulsion System
Japan Aerospace Exploration Agency (JAXA)
Drive technology
Propulsion System*
Health assessment
Link
60
PHM IEEE Data Challenge 2012 - FEMTO Bearing Data Set
FEMTO-ST Institute, Besançon, France
Mechanical component
Bearing
Prognosis
Link
61
PHM IEEE Data Challenge 2014 - Fuel Cell
FCLAB Research Federation (FR CNRS 3539, France)
Electrical component
Fuel Cell
Prognosis
Link
62
PHM IEEE Data Challenge 2023 - Planetary Gearbox
Mälardalen University (MDU) / Mälardalen Industrial Technology Center (MITC)
Mechanical component
Planetary Gear
Diagnosis
Link
63
PHM Society - Electromechanical Ball Screw Drive
GE Research and University of Tennessee Knoxville in part by the Advanced Research Projects Agency-Energy (ARPA-E), U.S. Department of Energy, under Award Number DE-AR0001290
Mechatronic system
Electromechanical Ball Screw Drive*
Diagnosis
Link
64
ResearchGate - Driveline Unbalanced Shaft
Giacomo Barbieri and David Sanchez-Londono and Laura Cattaneo and Luca Fumagalli and David Romero
Mechanical component
Unbalanced shaft (driveline)
Diagnosis
Link
65
SDOL - Diagnostics 101 bearing data
Kim, S., An, D., Choi, J-H. (2020). Diagnostics 101: A Tutorial for Fault Diagnostics of Rolling Element Bearing Using Envelope Analysis in MATLAB. Applied Sciences. 10(20):7302. doi:10.3390/app10207302
Mechanical component
Bearing
Fault detection
Link
66
SDOL - HS Gear
Creators Unknown
Mechanical component
Bearing
Fault detection
Link
67
SDOL - KAU Gear
Creators Unknown
Mechanical component
Bearing
Diagnosis
Link
68
Toyota Research Institute - Battery Cycle Life
Severson, K.A., Attia, P.M., Jin, N. et al. Data-driven prediction of battery cycle life before capacity degradation. Nat Energy 4, 383–391 (2019). https://doi.org/10.1038/s41560-019-0356-8
Electrical component
Battery
Prognosis
Link
69
UCI - AI4I 2020 Predictive Maintenance
Matzka, S. (2020). “Explainable Artificial Intelligence for Predictive Maintenance Applications," Proceedings of the Third International Conference on Artificial Intelligence for Industries (AI4I), pp. 69-74, doi: 10.1109/AI4I49448.2020.00023 .
Production system
Manufacturing Machine
Diagnosis
Link
70
UCI - Condition Based Maintenance of Naval Propulsion Plants Data Set
1: DIBRIS - University of Genoa 2: School of Marine Science and Technology, Newcastle University
Drive technology
Naval Propulsion Plant*
Prognosis
Link
71
UCI - Condition monitoring of hydraulic systems Data Set
ZeMA - Center for Mechatronics and Automation Technology Saarbrücken, Germany
Mechatronic system
Hydraulic System
Prognosis
Link
72
UCI - Mechanical Analysis
F. Bergadano, A. Giordana, L. Saitta, (1990). University of Torino, Italy
Mechatronic system
Electromechanical Devices
Diagnosis
Link
73
UCI - Robot Execution Failures Data Set
Universidade Nova de Lisboa, Monte da Caparica, Portugal
Robotic
Industrial Robot
Diagnosis
Link
74
UCI - Steel Plates Faults
Semeion Research Center of Sciences of Communication - Rome, Italy
Materials
Steel Plates Faults
Diagnosis
Link
75
Uni Lulea - Vibration Signals Wind Turbines
Luleå University of Technology: Martin del Campo Barraza, Sergio (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Sandin, Fredrik (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Strömbergsson, Daniel (Department of Engineering Sciences and Mathematics, Machine Elements)
Mechanical component
Bearing (Wind Turbines)
Fault detection
Link
76
Uni Paderborn KAt - Bearing Damage
University Paderborn Kat
Mechanical component
Bearing
Diagnosis
Link
77
Virkler
Virkler, D. A., Hillberry, B. M. and Goel, P. K. (1979). The Statistical Nature of Fatigue Crack Propagation. Journal of Engineering Materials and Technology, vol. 101, 148–153.
Materials
Aluminum Plate
Prognosis
Link
78
WASEDA - Fault Detection and Classification
De Bruijn, B., Nguyen, T. A., Bucur, D., & Tei, K. (2016). Benchmark datasets for fault detection and classification in sensor data. In A. Ahrens, O. Postolache, & C. Benavente-Peces (Eds.), SENSORNETS 2016 - Proceedings of the 5th International Confererence on Sensor Networks, pp. 185-195. . SciTePress. https://doi.org/10.5220/0005637901850195
Electrical component
Sensors
Diagnosis
Link
79
Zenodo - Predictive maintenance dataset
Huawei Munich Research Center
Mechatronic system
Elevator
Prognosis
Link

### v2

@@ 52 Oxford - Oxford Battery Degradation Dataset
Signals: current, temperature, voltage Prognosis Link
53 PHM Data Challenge 2008 - Turbofan
NASA Prognostics Center of Excellence Drive technology
Aircraft Engine*
Signals: anonymized, operating condition Prognosis Link
54 PHM Data Challenge 2009 - Gearbox Fault Detection
PHM Society, Gearbox fault detection data set, 2010 Mechanical component
Gear System (Gears, Bearings, Shafts)
Signals: speed, vibration Diagnosis Link
55 PHM Data Challenge 2010 - CNC milling machine cutters
X. Li, B.S. Lim, J.H. Zhou, S. Huang, S.J. Phua, K.C. Shaw, and M.J. Er; Singapore Institute of Manufacturing Technology, 71 Nanyang Drive, Singapore 638075 School of Electrical and Electronic Engineering, Nanyang Technological University, Nanyang Avenue, Singapore 639798 Manufacturing process
CNC Milling Machine Cutters
Signals: acoustic emission, force, vibration Prognosis Link
56 PHM Data Challenge 2011 - Anemometer Fault Detection
Creators Unknown Mechanical component
Anemometer
Signals: direction, speed, temperature Fault detection Link
57 PHM Data Challenge 2013 - Unknown
Creators Unknown Unknown
unknown
Signals: anonymized Health assessment Link
58 PHM Data Challenge 2014 - Unknown
Creators Unknown Unknown
unknown
Signals: anonymized Health assessment Link
59 PHM Data Challenge 2015 - Plant Fault Detection
Creators Unknown Production system
Industrial Plant Monitoring
Signals: anonymized Diagnosis Link
60 PHM Data Challenge 2016 - Semiconductor CMP
Crystec Technology Trading GmbH Manufacturing process
Chemicalmechanical Planarization System
Signals: flow rate, operating condition, pressure, speed Prognosis Link
61 PHM Data Challenge 2017 - Bogie Vehicle
Creators Unknown Mechanical component
Bogie
Signals: vibration Diagnosis Link
62 PHM Data Challenge 2018 - Ion Mill in Waver Manufactoring
Kai Goebel (at this time at NASA) Production system
Ion Mill Etching Tool
Signals: current, pressure, speed Prognosis Link
63 PHM Data Challenge 2019 - Fatigue Cracks
Creators Unknown Materials
Aluminum Structure with Dynamic Loading
Signals: vibration Prognosis Link
64 PHM Data Challenge 2020 Europe - Filtration System
Eker, O. F., Camci, F., & Jennions, I. K. (2016). Physics-based prognostic modelling of filter clogging phenomena. Mechanical Systems and Signal Processing, 75, pp. 395-412. Process technology
Filtration
Signals: flow rate, pressure Prognosis Link
65 PHM Data Challenge 2021 - Turbofan 2
M. Chao, C.Kulkarni, K. Goebel and O. Fink (2021). ”Aircraft Engine Runto-Failure Dataset under real flight conditions”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA (additional validation data compared to data set NASA - Turbofan engine degradation simulation data set 2 ) Drive technology
Aircraft Engine*
Signals: anonymized, operating condition Prognosis Link
66 PHM Data Challenge 2021 Europe: SCARA-robot
Swiss Centre for Electronics and Microtechnology (CSEM) Robotic
Industrial Robot
Signals: current, duration, heat slope, humidity, position, pressure, speed, temperature Diagnosis Link
67 PHM Data Challenge 2022 - Rock Drills
Epiroc Rock Drills AB, Örebro, 702 25, Sweden and Linköping University, Linköping, 581 83, Sweden Mechatronic system
Rock Drills
Signals: pressure Diagnosis Link
68 PHM Data Challenge 2022 Europe - PCB Production line
Danilo Giordano and Martino Trevisan, (2022). ”PHME Data Challenge”. European conference of the prognostics and health management society. / BITRON: Strada del Portone 95 - 10095 Grugliasco (TO), Italy, https://www.bitron.com/en Production system
PCB Production Line
Signals: inspection data Diagnosis Link
69 PHM Data Challenge 2023 - Gearbox
PHM Society, Gearbox data set, 2023 Mechanical component
Gear
Signals: speed, vibration Diagnosis Link
70 PHM Data Challenge 2023 Asia Pacific - Experimental Propulsion System
Japan Aerospace Exploration Agency (JAXA) Drive technology
Propulsion System*
Signals: pressure Health assessment Link
71 PHM Data Challenge 2024 - Helicopter Turbine Engines
Prognostics and Health Management (PHM) Society Drive technology
Helicopter Turbine Engine
Signals: power, speed, temperature, torque Fault detection Link
72 PHM IEEE Data Challenge 2012 - FEMTO Bearing Data Set
FEMTO-ST Institute, Besan¸con, France Mechanical component
@@ 72 PHM IEEE Data Challenge 2012 - FEMTO Bearing Data Set
Signals: temperature, vibration Prognosis Link
73 PHM IEEE Data Challenge 2014 - Fuel Cell
FCLAB Research Federation (FR CNRS 3539, France) Electrical component
Fuel Cell
Signals: current, current density, flow rate, hygrometry, pressure, temperature, voltage Prognosis Link
74 PHM IEEE Data Challenge 2023 - Planetary Gearbox
Mälardalen University (MDU) / Mälardalen Industrial Technology Center (MITC) Mechanical component
Planetary Gear
Signals: vibration Diagnosis Link
75 PHM Society - Electromechanical Ball Screw Drive
GE Research and University of Tennessee Knoxville in part by the Advanced Research Projects AgencyEnergy (ARPA-E), U.S. Department of Energy, under Award Number DEAR0001290 Mechatronic system
Electromechanical Ball Screw Drive*
Signals: current, position, speed, torque Diagnosis Link
76 ResearchGate - Driveline Unbalanced Shaft
Giacomo Barbieri and David SanchezLondono and Laura Cattaneo and Luca Fumagalli and David Romero Mechanical component
Unbalanced shaft (driveline)
Signals: vibration Diagnosis Link
77 SDOL - Diagnostics 101 bearing data
Kim, S., An, D., Choi, J-H. (2020). Diagnostics 101: A Tutorial for Fault Diagnostics of Rolling Element Bearing Using Envelope Analysis in MATLAB. Applied Sciences. 10(20):7302. doi:10.3390/app10207302 Mechanical component
Bearing
Signals: vibration Diagnosis Link
78 SDOL - HS Gear
Creators Unknown Mechanical component
Bearing
Signals: vibration Fault detection Link
79 SDOL - KAU Gear
Creators Unknown Mechanical component
Bearing
Signals: encoder Diagnosis Link
80 Toyota Research Institute - Battery Cycle Life
Severson, K.A., Attia, P.M., Jin, N. et al. Data-driven prediction of battery cycle life before capacity degradation. Nat Energy 4, 383–391 (2019). https://doi.org/10.1038/s41560-0190356-8 Electrical component
@@ 80 Toyota Research Institute - Battery Cycle Life
Signals: current, temperature, voltage Prognosis Link
81 UBFC - AMPERE Detection and diagnostics of rotor and stator faults in rotating machines
Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche-Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire Génie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère), doi: 10.25666/DATAUBFC-2023-03-06-03 Drive technology
Squirrel cage motor (Rotor, Stator, Bearing)
Signals: current, speed, vibration, voltage Diagnosis Link
82 UBFC - LASPI Detection and diagnostics of gearbox faults
Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche-Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire Génie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère), doi: 10.25666/DATAUBFC-2023-03-06 . Mechanical component
Bearing + Gear
Signals: current, vibration, voltage Diagnosis Link
83 UBFC - METALLICADOUR Detection and diagnostics of multi-axis robot faults
Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire G´enie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère). Robotic
Multi-axes robot (cutting tool + axes drifts)
Signals: current, force, position, torque, vibration Diagnosis Link
84 UCI - AI4I 2020 Predictive Maintenance
Matzka, S. (2020). “Explainable Artificial Intelligence for Predictive Maintenance Applications,” Proceedings of the Third International Conference on Artificial Intelligence for Industries (AI4I), pp. 69-74, doi: 10.1109/AI4I49448.2020.00023 . Production system
Manufacturing Machine
Signals: speed, temperature, torque, wear Diagnosis Link
85 UCI - Condition Based Maintenance of Naval Propulsion Plants Data Set
1: DIBRIS - University of Genoa 2: School of Marine Science and Technology, Newcastle University Drive technology
Naval Propulsion Plant*
Signals: flow rate, injection control, pressure, speed, temperature, torque Prognosis Link
86 UCI - Condition monitoring of hydraulic systems Data Set
ZeMA - Center for Mechatronics and Automation Technology Saarbrücken, Germany Mechatronic system
Hydraulic System
Signals: efficiency, flow rate, power, pressure, temperature, vibration Diagnosis Link
87 UCI - Mechanical Analysis 1990
F. Bergadano, A. Giordana, L. Saitta, (1990). University of Torino, Italy Mechatronic system
Electromechanical Devices
Signals: speed Diagnosis Link
88 UCI - Robot Execution Failures Data Set
Universidade Nova de Lisboa, Monte da Caparica, Portugal Robotic
Industrial Robot
Signals: force, torque Diagnosis Link
89 UCI - Steel Plates Faults
Semeion Research Center of Sciences of Communication - Rome, Italy Materials
Steel Plates Faults
Signals: geometry, inspection data, operating condition Diagnosis Link
90 Uni Lulea - Vibration Signals Wind Turbines
Lulea University of Technology: Martin del Campo Barraza, Sergio (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Sandin, Fredrik (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Strömbergsson, Daniel (Department of Engineering Sciences and Mathematics, Machine Elements) Mechanical component
Bearing (Wind Turbines)
Signals: vibration Fault detection Link
91 Uni Paderborn KAt - Bearing Damage
University Paderborn Kat Mechanical component
Bearing
Signals: speed, temperature, torque Diagnosis Link
92 Virkler - Fatigue Crack Propagation
Virkler, D. A., Hillberry, B. M. and Goel, P. K. (1979). The Statistical Nature of Fatigue Crack Propagation. Journal of Engineering Materials and Technology, vol. 101, 148–153. Materials
Aluminum Plate
Signals: crack length Prognosis Link
93 WASEDA - Fault Detection and Classification
De Bruijn, B., Nguyen, T. A., Bucur, D., & Tei, K. (2016). Benchmark datasets for fault detection and classification in sensor data. In A. Ahrens, O. Postolache, & C. Benavente-Peces (Eds.), SENSORNETS 2016 - Proceedings of the 5th International Confererence on Sensor Networks, pp. 185-195. . SciTePress. https://doi.org/10.5220/0005637901850195 Electrical component
Sensors
Signals: light, temperature Diagnosis Link
94 Zenodo - Ball bearings subjected to time-varying load and speed conditions
Aimiyekagbon, O. K. (2024). Run-tofailure data set of ball bearings subjected to time-varying load and speed conditions. Zenodo, doi: 10.5281/zenodo.10805043, University Paderborn Mechanical component
Bearing
Signals: temperature, vibration Prognosis Link
95 Zenodo - Data-driven capacity estimation of commercial lithium-ion batteries from voltage relaxation
Jiangong Zhu, Yixiu Wang, Yuan Huang, R. Bhushan Gopaluni, Yankai Cao, Michael Heere, Martin J. Mühlbauer, Liuda Mereacre, Haifeng Dai, Xinhua Liu, Anatoliy Senyshyn, Xuezhe Wei, Michael Knapp, & Helmut Ehrenberg. (2022). Data-driven capacity estimation of commercial lithium-ion batteries from voltage relaxation [Data set]. Zenodo. https://doi.org/10.5281/zenodo.6405084 Electrical component
@@ 95 Zenodo - Data-driven capacity estimation of commercial lithium-ion batteries from voltage relaxation
Signals: current, temperature, voltage Prognosis Link
96 Zenodo - Identifying degradation patterns of lithium ion batteries from impedance spectroscopy using machine learning
Zhang, Y., Tang, Q., Zhang, Y., Wang, J., Stimming, U., & Lee, A. A. (2020). Identifying degradation patterns of lithium ion batteries from impedance spectroscopy using machine learning [Data set]. Zenodo. https://doi.org/10.5281/zenodo.3633835 Electrical component
@@
Signals: current, temperature, voltage Prognosis Link
97 Zenodo - Predictive maintenance dataset
Huawei Munich Research Center Mechatronic system
Elevator
Signals: humidity, vibration Prognosis Link
98 Zenodo - TBSI Sunwoda Battery Dataset
Shengyu, T. (2024). TBSI Sunwoda Battery Dataset [Data set]. Zenodo. https://doi.org/10.5281/zenodo.10715209 Electrical component
@@ 98 Zenodo - TBSI Sunwoda Battery Dataset
Signals: current, voltage Prognosis Link